Linux HOWTO Index

by Matt Welsh, mdw@sunsite.unc.edu v2.0, 17 March 1994

This is an index to the Linux HOWTO documents.

Contents

1 What are Linux HOWTOs? 1

2 Where to get Linux HOWTOs 1

3 HOWTO Index 2

4 Wanted HOWTOs 6

5 Writing and submitting a HOWTO 6

6 Miscellaneous and Legalese 7

1 What are Linux HOWTOs?

Linux HOWTOs are short online documents which describe in detail a certain aspect of configuring or using the Linux system. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET-2-HOWTO (previously the NET-2-FAQ) and the Printing HOWTO.

Information in HOWTOs is generally more detailed and in-depth than what can be squeezed into the Linux FAQ. For this reason, the Linux FAQ is being rewritten. A large amount of the information contained therein will be relegated to various HOWTO documents. The FAQ will be a shorter list of frequently asked questions about Linux, covering small specific topics. Most of the “useful” information in the FAQ will now be covered in the HOWTOs.

HOWTOs are comprehensive docs—much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. For example, the NET-2-FAQ has been renamed to the NET-2-HOWTO, because it wasn’t in question-and-answer format. However, you will see the NET-2-HOWTO named as the NET-2-FAQ in many places. The two docs are one and the same.

2 Where to get Linux HOWTOs

HOWTOs can be retrieved via anonymous FTP from the following sites:

• sunsite.unc.edu:/pub/Linux/docs/HOWTO
• tsx-11.mit.edu:/pub/linux/docs/HOWTO

as well as the many mirror sites, which are listed in the Linux META-FAQ (see below).

The Index, printed below, lists the currently available HOWTOs.

HOWTOs are also posted regularly to the newsgroups comp.os.linux and comp.os.linux.announce. In addition, a number of the HOWTOs will be crossposted to news.answers. Therefore, you can find the Linux

HOWTOs on the news.answers archive site rtfm.mit.edu.

3 HOWTO Index

The following Linux HOWTOs are currently available.

• Linux Bootdisk HOWTO, by Graham Chapman <grahamc@zeta.org.au>. How to create a boot/root maintenance disk for Linux

• Linux Busmouse HOWTO, by mike@starbug.apana.org.au (Mike Battersby). Information on bus mouse compatibility with Linux

• Linux CDROM HOWTO, by tranter@software.mitel.com (Jeff Tranter). Information on CD-ROM drive compatibility for Linux

• Linux Commercial HOWTO, by hm@ix.de (Harald Milz). Listing of commercial software prducts for Linux

• Linux DOSEMU HOWTO, by deisher@enws125.EAS.ASU.EDU (Michael E. Deisher). HOWTO about the Linux MS-DOS Emulator, DOSEMU

• Linux Distribution HOWTO, by mdw@sunsite.unc.edu (Matt Welsh). A list of mail order distributions and other commercial services

• Linux Ethernet HOWTO, by Paul Gortmaker <gpg109@rsphysse.anu.edu.au>. Information on Ethernet hardware compatibility for Linux

• Linux Ftape HOWTO, by khp@pip.dknet.dk (Ftape-HOWTO maintainer). Information on ftape drive compatibility with Linux

• Linux German HOWTO, by Winfried Truemper <truemper@MI.Uni-Koeln.DE>. Information on using Linux with German-specific features

• Linux HAM HOWTO, by terryd@extro.ucc.su.oz.au (Terry Dawson). HOWTO configure amateur radio software for Linux

• Linux HOWTO Index, by mdw@sunsite.unc.edu (Matt Welsh). Index of HOWTO documents about Linux.

• Linux Hardware Compatibility HOWTO, by frac@ksc.au.ac.th (Tawei Wan). A list of hardware known to work with Linux

• Linux Installation HOWTO, by mdw@sunsite.unc.edu (Matt Welsh). How to obtain and install the Linux software

• Linux JE-HOWTO, by Yasuhiro Yamazaki <hiro@rainbow.physics.utoronto.ca>. Information on JE, a set of Japanese language extensions for Linux

• Linux Kernel HOWTO, by Brian Ward <ward@blah.tu-graz.ac.at>. Upgrading and compiling the Linux kernel

• Linux Keystroke HOWTO, by Zenon Fortuna <zenon@netcom.com>. HOWTO bind macro actions to keystrokes under Linux

• Linux MGR HOWTO, by broman@Np.nosc.mil (Vincent Broman). Information on the MGR graphics interface for Linux

• Linux Electronic Mail HOWTO, by vince@victrola.wa.com (Vince Skahan). Information on Linuxbased mail servers and clients

• Linux NET-2 HOWTO, by terryd@extro.ucc.su.oz.au (Terry Dawson). HOWTO configure TCP/IP networking, SLIP, PLIP, and PPP under Linux.

• Linux NIS HOWTO, by erwin@nioz.nl (Erwin Embsen). Information on using NIS/YP on Linux systems

• Linux News HOWTO, by vince@victrola.wa.com (Vince Skahan). Information on USENET news server and client software for Linux

• Linux PCI-HOWTO, by Michael Will <michaelw@desaster.student.uni-tuebingen.de>. Information on PCI-architecture compatibility with Linux

• Linux PPP-HOWTO, by Al Longyear <longyear@netcom.com>. Information on using PPP networking with Linux

• Linux Printing HOWTO, by gtaylor@cs.tufts.edu (Grant Taylor). HOWTO on printing software for Linux

• Linux SCSI HOWTO, by Drew Eckhardt <drew@kinglear.cs.Colorado.EDU>. Information on SCSI driver compatibility with Linux

• Linux Serial HOWTO, by gregh@cc.gatech.edu (Greg Hankins). Information on use of serial devices and communications software

• Linux Sound HOWTO, by tranter@software.mitel.com (Jeff Tranter). Sound hardware and software for the Linux operating system

• Linux Term HOWTO, by Patrick Reijnen (patrickr@cs.kun.nl). HOWTO use the ‘term’ communications package on Linux systems • Linux Tips HOWTO, by Vince Reed <reedv@rpi.edu>. HOWTO on miscellaneous tips and tricks

for Linux

• Linux UPS HOWTO, by Harvey J. Stein <hjstein@math.huji.ac.il>. Information on using a UPS power supply with Linux

• Linux UUCP HOWTO, by vince@victrola.wa.com (Vince Skahan). Information on UUCP software for Linux

• Linux XFree86 HOWTO, by geyer@polyhymnia.iwr.uni-heidelberg.de (Helmut Geyer). HOWTO on installation of XFree86 (X11R5) for Linux

• Linux Bootdisk HOWTO, by Graham Chapman <grahamc@zeta.org.au>. How to create a boot/root maintenance disk for Linux

• Linux Busmouse HOWTO, by mike@starbug.apana.org.au (Mike Battersby). Information on bus mouse compatibility with Linux

• Linux CDROM HOWTO, by tranter@software.mitel.com (Jeff Tranter). Information on CD-ROM drive compatibility for Linux

• Linux Commercial HOWTO, by hm@ix.de (Harald Milz). Listing of commercial software prducts for Linux

• Linux DOSEMU HOWTO, by deisher@enws125.EAS.ASU.EDU (Michael E. Deisher). HOWTO about the Linux MS-DOS Emulator, DOSEMU

• Linux Distribution HOWTO, by mdw@sunsite.unc.edu (Matt Welsh). A list of mail order distributions and other commercial services

• Linux Ethernet HOWTO, by Paul Gortmaker <gpg109@rsphysse.anu.edu.au>. Information on Ethernet hardware compatibility for Linux

• Linux Ftape HOWTO, by khp@pip.dknet.dk (Ftape-HOWTO maintainer). Information on ftape drive compatibility with Linux

• Linux German HOWTO, by Winfried Truemper <truemper@MI.Uni-Koeln.DE>. Information on using Linux with German-specific features

• Linux HAM HOWTO, by terryd@extro.ucc.su.oz.au (Terry Dawson). HOWTO configure amateur radio software for Linux

• Linux HOWTO Index, by mdw@sunsite.unc.edu (Matt Welsh). Index of HOWTO documents about Linux.

• Linux Hardware Compatibility HOWTO, by frac@ksc.au.ac.th (Tawei Wan). A list of hardware known to work with Linux

• Linux INFO-SHEET, by johnsonm@sunsite.unc.edu (Michael K. Johnson). Generic introduction to the Linux operating system • Linux Installation HOWTO, by mdw@sunsite.unc.edu (Matt Welsh). How to obtain and install

the Linux software

• Linux JE-HOWTO, by Yasuhiro Yamazaki <hiro@rainbow.physics.utoronto.ca>. Information on JE, a set of Japanese language extensions for Linux

• Linux Kernel HOWTO, by Brian Ward <ward@blah.tu-graz.ac.at>. Upgrading and compiling the Linux kernel

• Linux Keystroke HOWTO, by Zenon Fortuna <zenon@netcom.com>. HOWTO bind macro actions to keystrokes under Linux

• Linux META-FAQ, by johnsonm@sunsite.unc.edu (Michael K. Johnson). A listing of Linux sources of information

• Linux MGR HOWTO, by broman@Np.nosc.mil (Vincent Broman). Information on the MGR graphics interface for Linux

• Linux Electronic Mail HOWTO, by vince@halcyon.com (Vince Skahan). Information on Linuxbased mail servers and clients

• Linux NET-2 HOWTO, by terryd@extro.ucc.su.oz.au (Terry Dawson). HOWTO configure TCP/IP networking, SLIP, PLIP, and PPP under Linux.

• Linux NIS HOWTO, by erwin@nioz.nl (Erwin Embsen). Information on using NIS/YP on Linux systems

• Linux News HOWTO, by vince@halcyon.com (Vince Skahan). Information on USENET news server and client software for Linux

• Linux PCI-HOWTO, by Michael Will <michaelw@desaster.student.uni-tuebingen.de>. Information on PCI-architecture compatibility with Linux

• Linux PPP-HOWTO, by Al Longyear <longyear@netcom.com>. Information on using PPP networking with Linux

• Linux Printing HOWTO, by gtaylor@cs.tufts.edu (Grant Taylor). HOWTO on printing software for Linux

• Linux SCSI HOWTO, by Drew Eckhardt <drew@kinglear.cs.Colorado.EDU>. Information on SCSI driver compatibility with Linux

• Linux SCSI Programming HOWTO, by Heiko Eissfeldt <heiko@colossus.escape.de>. Information on programming the generic Linux SCSI interface

• Linux Serial HOWTO, by gregh@cc.gatech.edu (Greg Hankins). Information on use of serial devices and communications software

• Linux Sound HOWTO, by tranter@software.mitel.com (Jeff Tranter). Sound hardware and software for the Linux operating system • Linux Term HOWTO, by Patrick Reijnen (patrickr@cs.kun.nl). HOWTO use the ‘term’ communi-

cations package on Linux systems

• Linux Tips HOWTO, by Vince Reed <reedv@rpi.edu>. HOWTO on miscellaneous tips and tricks for Linux

• Linux UPS HOWTO, by Harvey J. Stein <hjstein@math.huji.ac.il>. Information on using a UPS power supply with Linux

• Linux UUCP HOWTO, by vince@halcyon.com (Vince Skahan). Information on UUCP software for Linux

• Linux XFree86 HOWTO, by geyer@polyhymnia.iwr.uni-heidelberg.de (Helmut Geyer). HOWTO on installation of XFree86 (X11R5) for Linux

4 Wanted HOWTOs

I would be very happy if someone would write the following HOWTO documents, and submit them to me (see the instructions in the next section below). Most of these are modifications of existing FAQ documents, and should be easy to put into HOWTO format. Keep in mind that HOWTOs generally aren’t in FAQ format, but can contain an FAQ section at the end.

Some of these are under construction; however, if you’re interested in working on one of them, please let me know.

• GCC/LIBRARY HOWTO. Discussing installation/upgrading of the GCC and shared libraries under Linux. How to build shared libraries. Most of the material would come from Mitchum DeSouza’s GCC FAQ.

• If you have an idea for another HOWTO, please mail me!

5 Writing and submitting a HOWTO

If you would like to write a Linux HOWTO document, there are a few guidelines that you should follow.

• Format the document neatly. HOWTOs must be available in plain ASCII format, but you are free to use a formatting tool (texinfo, LameTeX, nroff) to format the document. Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• You may wish to use the Linuxdoc-SGML package, available from ftp.cs.cornell.edu:/pub/mdw/, to format the HOWTO. This package allows you to produce LaTeX, plain ASCII, and HTML from a single source document, and was designed specifically with the HOWTOs in mind. If you are writing a new HOWTO you should seriously consider this.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname).

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an ”FAQ” or ”Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs as a model.

After you have written the HOWTO, mail it to me (Matt Welsh) at mdw@sunsite.unc.edu. If you have use Linuxdoc-SGML, simply mail me the SGML source; I take care of formatting the documents. I’ll also take care of archiving the HOWTO and posting it to the various newsgroups.

It is important that you go through me when submitting a HOWTO, as I maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

All you have to do is send me periodic updates (every month or so, as needed) and I will take care of posting the HOWTO regularly.

6 Miscellaneous and Legalese

If you have questions, please feel free to mail mdw@sunsite.unc.edu. The Linux FAQ rewrite is being coordinated by Ian Jackson, ijackson@nyx.cs.du.edu, with help from others.

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, without permission of the author. Translations and derivative works are similarly permitted without express permission. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs. If you have questions, please contact Matt Welsh, the Linux HOWTO

coordinator, at mdw@sunsite.unc.edu.